# Optimal approximation of infinite-dimensional holomorphic functions II: recovery from i.i.d. pointwise samples

### v2

@@ Appendix A Supporting results on Jacobi polynomials
In the following result, we show the previously-made claim that sequences $\bm{\alpha},\bm{\beta}$ satisfying Assumption 4.1 also satisfy Assumption 3.
Proposition A.1.
Suppose that the sequences $\bm{\alpha},\bm{\beta}$ satisfy Assumption 4.1. Then the measure (4.3) satisfies Assumption 3.
Proof.
Let $\sigma_{i}$ and $\tau_{i}$ be as in (3.5). Then the function $\frac{y-\tau_{i}}{\sigma_{i}}$ is precisely the order-one, orthonormal Jacobi polynomial $\Psi^{\alpha,\beta}_{1}(y)$ with $\alpha=\alpha_{i}$, $\beta=\beta_{i}$. Since (see, e.g., [1, §18.5.5])
$P^{\alpha,\beta}_{1}(y)=\frac{1}{2}(\alpha-\beta)+\frac{\alpha+\beta+2}{2}y,$
(A.2) implies that
$\frac{1}{\sigma_{i}}=\left(\frac{h^{\alpha,\beta}_{0}}{h^{\alpha,\beta}_{1}}\right)^{1/2}\left(\frac{\alpha+\beta+2}{2}\right).$
Using (A.1) and the fact that $z\Gamma(z)=\Gamma(z+1)$ for $z>0$, we get
$\frac{h^{\alpha,\beta}_{0}}{h^{\alpha,\beta}_{1}}=\frac{\alpha+\beta+3}{(\alpha+1)(\beta+1)},$ (A.3)
and therefore
$\sigma_{i}=\frac{2}{\alpha+\beta+2}\sqrt{\frac{(\alpha+1)(\beta+1)}{\alpha+\beta+3}}.$
We now apply Assumption (4.1), to deduce that
$\sigma_{i}\geq\frac{1}{1/\tau+1}\frac{\tau}{\sqrt{2/\tau+3}}.$
Since the right-hand side is independent of $i$, the result follows. ∎
Next, we need to show a key technical lemma (Lemma A.3) concerning the uniform growth of the Jacobi polynomials. This lemma was used in the proofs of Theorems 5.15.2. Its proof relies on the following result.
@@
Finally, to end this section we consider the following assumption, which we show is equivalent to Assumption 4.1.
@@
${\|\Psi^{\alpha,\beta}_{\nu}\|}^{2}_{L^{\infty}([-1,1])}\geq\max\{|\Psi^{\alpha,\beta}_{\nu}(\pm 1)|^{2}\}=\left(\frac{h^{\alpha,\beta}_{0}}{h^{\alpha,\beta}_{\nu}}\right)\max\left\{{\nu+\alpha\choose\nu},{\nu+\beta\choose\nu}\right\}.$
@@ Appendix B The relationship between $\overline{\theta_{m}}(p,\mathsf{M})$ and $\overline{\theta_{m}}(p)$
Let $\varrho$ be a tensor-product probability measure as in (2.1), $\bm{b}=(b_{i})_{i\in\mathbb{N}}\in[0,\infty)^{\mathbb{N}}$ and $\pi:\mathbb{N}\rightarrow\mathbb{N}$ be a bijection. Consider the quantity $\theta_{m}(\bm{b})$ defined in (3.2). Then
@@
where $\Theta_{m}(\cdot;\cdot,\cdot)$ is as in (3.1), $\bm{b}_{\pi}=(b_{\pi(i)})_{i\in\mathbb{N}}$ and $\varrho_{\pi}=\varrho_{\pi(1)}\times\varrho_{\pi(2)}\times\cdots$ is the permuted measure. In particular,
$\theta_{m}(\bm{b})=\theta_{m}(\bm{b}_{\pi})$
whenever $\varrho$ is a tensor product of the same measure, i.e., $\varrho_{1}=\varrho_{2}=\cdots$ in (2.1).
@@
The first equality is simply the definition (3.2) of $\theta_{m}(\bm{b})$. We now prove the second equality. By (3.1),
@@
The previous lemma implies that $\theta_{m}(\bm{b})$ generally does not coincide with $\theta_{m}(\bm{b}_{\pi})$, due to the change of measure from $\varrho$ to $\varrho_{\pi}$. However, this does hold whenever $\varrho$ is a tensor product of the same measure. We now show in this case that the terms $\overline{\theta_{m}}(p)$ and $\overline{\theta_{m}}(p,\mathsf{M})$ in (3.3) also coincide whenever this holds.
@@
Let $\varrho$ be a tensor-product probability measure as in (2.1) for which $\varrho_{1}=\varrho_{2}=\cdots$. Then the terms in (3.3) coincide, i.e., $\overline{\theta_{m}}(p)=\overline{\theta_{m}}(p,\mathsf{M})$.
@@ Appendix C Notation
This table summarizes main notation used in the paper. Note that in the paper we also consider scalar-valued versions of various spaces and norms, in which case we drop the $;\mathcal{V}$ symbol used below.

### v3

@@ Appendix A Supporting results on Jacobi polynomials
We now show a key technical lemma (Lemma A.2) on the uniform growth of the Jacobi polynomials. This lemma was used in the proofs of Theorems 5.25.3. Its proof relies on the following result.
@@
Next, we consider the following assumption, which we show is equivalent to Assumption 4.1 (recall Remark 5.1).
@@
${\|\Psi^{\alpha,\beta}_{\nu}\|}^{2}_{L^{\infty}([-1,1])}\geq\max\{|\Psi^{\alpha,\beta}_{\nu}(\pm 1)|^{2}\}=\left(\frac{h^{\alpha,\beta}_{0}}{h^{\alpha,\beta}_{\nu}}\right)\max\left\{{\nu+\alpha\choose\nu},{\nu+\beta\choose\nu}\right\}^{2}.$
@@ Appendix B The relationship between $\overline{\theta_{m}}(p,\mathsf{M})$ and $\overline{\theta_{m}}(p)$
Let $\varrho=\varrho_{1}\times\varrho_{2}\times\cdots$ be a tensor-product probability measure on $\mathcal{U}$, where each $\varrho_{i}$ is a probability measure on $[-1,1]$, $\bm{b}=(b_{i})_{i\in\mathbb{N}}\in[0,\infty)^{\mathbb{N}}$ and $\pi:\mathbb{N}\rightarrow\mathbb{N}$ be a bijection. Consider the quantity $\theta_{m}(\bm{b})$ defined in (3.2). Then
@@
where $\Theta_{m}(\cdot;\cdot,\cdot)$ is as in (3.1), $\bm{b}_{\pi}=(b_{\pi(i)})_{i\in\mathbb{N}}$ and $\varrho_{\pi}=\varrho_{\pi(1)}\times\varrho_{\pi(2)}\times\cdots$ is the permuted measure. In particular, if $\varrho_{1}=\varrho_{2}=\cdots$ then
$\theta_{m}(\bm{b})=\theta_{m}(\bm{b}_{\pi}).$
@@
The first equality is just the definition (3.2) of $\theta_{m}(\bm{b})$. We now prove the second. By (3.1),
@@
The previous lemma implies that $\theta_{m}(\bm{b})$ has no reason to coincide with $\theta_{m}(\bm{b}_{\pi})$, due to the change of measure from $\varrho$ to $\varrho_{\pi}$. However, this does hold whenever $\varrho$ is a tensor product of the same measure. We now show that the terms $\overline{\theta_{m}}(p)$ and $\overline{\theta_{m}}(p,\mathsf{M})$ in (3.3) also coincide whenever this holds.
@@
Let $\varrho=\varrho_{1}\times\varrho_{1}\times\cdots$ be a tensor-product probability measure on $\mathcal{U}$, where $\varrho_{1}$ is a probability measure on $[-1,1]$. Then the terms in (3.3) coincide, i.e., $\overline{\theta_{m}}(p)=\overline{\theta_{m}}(p,\mathsf{M})$.
@@ Appendix C Notation
Table 2 summarizes main notation used in the paper. Note that we also consider scalar-valued versions of various spaces and norms, in which case we drop the $;\mathcal{V}$ symbol used in the table.
Table 1: Table of notation.